Discuss with colleagues and IITians

in one dimensional collision between two identical particles A and B, B is stationary and A ha momentum p before impact. during impact b gives an impulse j to A. find coefficient of restitution.

somy chaturvedi
13 Years agoGrade

1 Answer

Akash Kumar Dutta
13 Years ago

Dear Somy,
relaive velocity of approach=p/m
loss in momentum for A= p-j
relative velocit of separation=  j/m - (p-j)/m= 2j - p/m
e=  2j-p/p = 2.(j/p) - 1
Regards.